Molecule 1 Molecule O

Show the synthesis from acetylene for K (C11 H14 O) and O (C5 H12 O) using the spectra below

Molecule K ЗН 3H 5H 2H diastereotopic 1н PPM oMolecule O 6H 4H 1H 1H exchangeable PPM
Answer #1

I (cllH14 o) = DBE = 11-(14)+1 :12-7=5 =ut1 - 83.1 (sertet, 14 1 Anma : - c4 c - 42 - sa-Icsi 34) 81.2(d, 3H) 51.5-2.0(ld ,2u